OpenStudy (anonymous):

Part 1: explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences. Part 2: If you were an Algebra 1 instructor and were creating a test on factoring trinomials of the form ax2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (anonymous):

\[2x ^{2}+13x+15\] 2*15=30 10*3=30 10+3=13 when multiplication of a & c is positive then break b in two parts say p & q then p+q=b & p*q=a*c